OpenStudy (anonymous):

Christa buys 14 flowerpots. Some are small and cost $2.40 each. The rest are large and cost $5.60 each. She spends a total of $49.60. Which equation models this situation? Let s represent the number of small flowerpots she buys.

OpenStudy (anonymous):

can someone just put it in question form so i can solve it

OpenStudy (anonymous):

Wow, Soo vaig, How many Small pots ??

OpenStudy (anonymous):

ill say can u help please

OpenStudy (anonymous):

Ill try, Tell me

OpenStudy (anonymous):

tell u what

OpenStudy (anonymous):

how many small flower pots ?

OpenStudy (anonymous):

idk @ZairaHernandez

OpenStudy (zairahernandez):

Let s be the number of small pots Let l be the number of large pots TotalCost=2.4s+5.6l 49.60=2.4s+5.6l Rearrange to get s on its own

OpenStudy (anonymous):

this is the entire question Christa buys 14 flowerpots. Some are small and cost $2.40 each. The rest are large and cost $5.60 each. She spends a total of $49.60. Which equation models this situation? Let s represent the number of small flowerpots she buys. @ZairaHernandez

OpenStudy (anonymous):

Wow, No help at all @ZairaHernandez

OpenStudy (anonymous):

Christa buys 14 flowerpots. Some are small and cost $2.40 each. The rest are large and cost $5.60 each. She spends a total of $49.60. Which equation models this situation? Let s represent the number of small flowerpots she buys. A. 5.6s + 2.4(s + 14) = 49.6 B. (2.4 + 9.6)(14 – s) = 49.6 C. 2.4s + 5.6(14 – s) = 49.6 D. 5.6s + 2.8(14) = 49.6

OpenStudy (anonymous):

@HamdyJS

OpenStudy (anonymous):

Wow Townie

OpenStudy (anonymous):

@linn99123

OpenStudy (anonymous):

@No.name